Whats wrong with this?
|m+3| = 12-2m
m+3=12-2m
3=12+2m
3=12-12+2m
3/2=2m/2
m=1.5 or -5


also
I need help with this one.
1/2|6-2x|=3x + 1

Answer by stanbon(75887)

|m+3| = 12-2m
m+3 = 12-2m or m+3=-(12-2m)
3m = 9 or m+3 =2m-12
m = 3 or m=15
================
(1/2)|6-2x|=3x + 1
Multiply both sides by 2 to get:
|6-2x|=6x+2
6-2x=6x+2 or 6-2x=-(6x+2)
8x=4 or 6-2x=-6x-2
x=(1/2) or 4x=-8
x=(1/2) or x=-2
==================
Cheers,
Stan H.